m/s2 for 10.0 s, then slows down to a stop with a constar SeA len sioning tomrest moves in a straigh line with a constant acceleration of 2.20 m/s for 100 s,then slows down to sp ith a consta of 3.05 m/s. How far does it travel? x 10.13 meters
Answer #1

Distance traveled while acceleration

= 1/2at2 ​​​​=1/2*2.20*10*10=110m

Distance travelled while stopping= s=v2/2a

Where v is 22 m/s

S= 22*22/2*3.05 =79.34

Total distance=110+79.34=189.34